OpenStudy (anonymous):

can someone help me, giving medals. The area of a rectangular piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. (5 points) Part B: What is the width of the land in meters? Show the steps of your work. (5 points)

OpenStudy (anonymous):

Ur square l n b will be like this (l-5) (w+1)

OpenStudy (anonymous):

? im confused

OpenStudy (anonymous):

can you help me write the equation and then i will solve it and show you all my work.

OpenStudy (anonymous):

l*w=280, w=280/l (l-5)(w+1)=l^2

OpenStudy (anonymous):

@dominiquebee72

OpenStudy (anonymous):

ok so is the second line an equation?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

(l-5)(280+1)=l^2 wait does L equal, or is that what im supposed to figure out?

OpenStudy (anonymous):

@Probhat

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

A solution using Mathematica 9 is attached.

OpenStudy (anonymous):

1st eq. l*w=280 2nd eq. 4(l-5+w+1)=4l @dominiquebee72

OpenStudy (anonymous):

for the second i got w=4

OpenStudy (anonymous):

the first one i got L=280/w

OpenStudy (anonymous):

@probhat